# PriceFM: Foundation Model for Probabilistic Electricity Price Forecasting

### v1

@@ Abstract
Electricity price forecasting in Europe presents unique challenges due to the continent’s increasingly integrated and physically interconnected power market. While recent advances in deep learning and foundation models have led to substantial improvements in general time series forecasting, most existing approaches fail to capture the complex spatial interdependencies and uncertainty inherent in electricity markets. In this paper, we address these limitations by introducing a comprehensive and up-to-date dataset across 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01. Building on this groundwork, we propose PriceFM, a spatiotemporal foundation model that integrates graph-based inductive biases to capture spatial interdependencies across interconnected electricity markets. The model is designed for multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting. Extensive experiments and ablation studies confirm the model’s effectiveness, consistently outperforming competitive baselines and highlighting the importance of spatial context in electricity markets. The dataset and code can be found at https://github.com/runyao-yu/PriceFM.
Introduction
@@
The European electricity market is physically interconnected through a network of cross-border transmission lines, enabling the exchange of electricity between regions and optimizing the social welfare at the European level (Lago et al. 2018). However, physical constraints, such as limited transmission capacity, can restrict electricity flow between regions and lead to zonal price differences (Finck 2021), illustrated in Figure 1. These price disparities highlight the spatial nature of electricity price formation. Recent studies show that electricity price dynamics are strongly influenced by spatial interdependencies and cannot be accurately captured using region-specific models (Do et al. 2024). Therefore, explicitly modeling the spatial structure of the European electricity market is essential for producing accurate price forecasts.
@@
Figure 1: Spatial distribution of electricity price and number of neighboring regions. a Electricity prices for 38 European regions averaged from 2022-01-01 to 2025-01-01. A significant zonal price difference is observed between north and south regions. b Number of neighboring regions that are directly connected to certain region via transmission lines. For example, France (FR) and Portugal (PT) are directly connected to Spain (ES), thus the number of regions for ES is 2.
@@
Most existing studies on electricity price forecasting do not explicitly model the spatial structure and focus on a single-region market, particularly Germany (Muniain and Ziel 2020; Maciejowska, Nitka, and Weron 2021; Kitsatoglou et al. 2024), as the German market is one of the largest markets in Europe. Other studies explore forecasting methods for markets such as Denmark, Finland, and Spain, also using region-specific models (Ziel and Weron 2018; Gianfreda, Ravazzolo, and Rossini 2020; Loizidis, Kyprianou, and Georghiou 2024). More recent works explicitly model the spatial nature of the electricity price. For instance, a Graph Convolutional Network (GCN) is applied to capture spatial interdependencies in the Nordic markets, such as Norway, Sweden, and Finland (Yang et al. 2024). Moreover, an attention-based variant is developed to predict prices in certain European markets such as Austria, Germany, and Hungary (Meng et al. 2024). However, these models cover only subsets of Europe and primarily produce point forecasts, failing to capture the uncertainty inherent in electricity prices.
@@
Uncertainty modeling in electricity markets is critical, as the electricity price is strongly influenced by intermittent renewable generation and fluctuating demand (Lago et al. 2021). Consequently, price forecasting should extend beyond traditional pointwise forecasting to explicitly quantify uncertainties, especially for applications involving risk-sensitive applications such as energy trading and operational planning (Ziel and Steinert 2018). A comprehensive survey summarizes various probabilistic forecasting approaches based on quantile regression methods (Lago et al. 2021). However, these existing methods often focus on single-region markets, thereby neglecting the rich spatial information.
@@
In recent years, foundation models for time series forecasting have achieved remarkable success across diverse domains, demonstrating strong generalization capabilities by capturing complex data patterns (Zhou et al. 2022; Wu et al. 2023; Wang et al. 2024), which makes them attractive candidates for electricity price forecasting. However, existing models are primarily designed for general univariate or multivariate time series tasks and are not explicitly tailored to electricity markets (Liang et al. 2024). Moreover, some foundation models only provide pointwise forecasts and thus fail to capture the uncertainty essential for risk-aware trading decisions in volatile markets. Therefore, addressing the spatial interdependencies and uncertainty in electricity price forecasting within the integrated European market necessitates a tailored foundation model.
@@
To support the development of foundation models for electricity price forecasting, there is a pressing need for high-quality, large-scale, and up-to-date datasets that reflect the spatiotemporal complexity of the integrated European market. However, existing datasets are often fragmented in structure, cover only short time periods, are outdated, or focus on individual regions (Lago et al. 2021). This lack of standardized, cross-border data poses a significant barrier to training and evaluating foundation models.
In this paper, we bridge the gap between recent advances in foundation models and the requirements of electricity price forecasting in spatially interconnected markets. To this end, we introduce a comprehensive and up-to-date dataset capturing key spatiotemporal drivers of electricity prices across Europe. As a step forward, we propose PriceFM, a foundation model that incorporates graph-based inductive biases to learn spatial structures and generate probabilistic forecasts tailored to the integrated European electricity market.
@@
  • We introduce and release a comprehensive, up-to-date dataset. To the best of our knowledge, this is the largest and most diverse open dataset for European electricity markets, comprising day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore), covering 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01.
  • We propose and release the PriceFM, a novel forecasting framework that integrates prior graph knowledge derived from the spatial topology of the European electricity market. PriceFM supports multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting.
  • We conduct extensive experiments to evaluate the model’s performance against multiple baselines, and assess the impact of design choices through ablation studies, thereby providing both quantitative evidence of overall performance and empirical insights into optimal configurations.
Preliminary
The forecasting target is $\mathcal{T}=24$ hourly prices for the delivery day $\mathcal{D}+1$, using data available before gate closure, typically around midday on day $\mathcal{D}$. After midday on $\mathcal{D}$, the electricity prices for $\mathcal{D}+1$ are published and known. We employ a backward-looking window of size L{L} (e.g. L=24{L}=24 corresponds to 24 hours from $\mathcal{D}$), for known electricity prices, denoted as $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$. We also include forward-looking exogenous features, such as day‑ahead forecasts of load, solar, and wind (onshore and offshore) power generation for $\mathcal{D}+1$, denoted as $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, made on $\mathcal{D}$ before gate closure, as well as their historical values over L{L}. The forecasting setup and the choice of feature set are widely used in prior works (Maciejowska 2020; Uniejewski and Weron 2021; Meng et al. 2024), illustrated in Figure 2. Note that this work aims to utilize multi-region inputs to produce multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasts. Therefore, the input and output are defined as:
@@
where $r_{\mathrm{in}},r_{\mathrm{out}}\in\mathcal{R}=\{\text{AT},\dots,\text{SK}\}$ (region codes, detailed in Appendix, Table 1.), $\tau\in\mathcal{Q}=\{0.1,0.5,0.9\}$ (quantile levels), $f_{1}=1$, and $f_{2}$ varies by region. For example, the forecasted wind power generation is not available in SI.
Figure 2: Forecasting setup and feature set.
Figure 3: European-level energy data in 2024, averaged across all available regions (depending on data availability). a Electricity price. Price spikes sharply during the morning and evening peak, dip around midday, and shows higher volatility in winter. b Forecasted load. Load exhibits a clear double‐peak each day (in morning and evening) with winter peaks substantially larger than summer. c Forecasted solar power generation. Solar is zero overnight, rises in a smooth bell curve to a strong midday maximum, then falls back to zero by dusk, and is much higher in summer. d Forecasted wind power generation (sum of onshore and offshore). Wind lacks a daily pattern, fluctuates with high‐frequency spikes, and is much higher in the windy winter.
Data
The dataset is primarily collected from the European Network of Transmission System Operators for Electricity (ENTSO-E). We release the raw and cleaned dataset.
Spatiotemporal Coverage
Spatially, the dataset covers 24 European countries (38 regions). These regions reflect transmission zones rather than administrative boundaries. For example, DK is split into two regions: DK1 and DK2. Each is connected to different neighboring regions, resulting in distinct cross-border power flows. Temporally, the dataset spans from 2022-01-01 to 2025-01-01, providing wide temporal coverage.
Feature Set
The feature set includes day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore). For simplicity, we refer to these features as price, load, solar, and wind (onshore and offshore), respectively. Importantly, some regions only contain partial features. The full availability of features across regions is detailed in Appendix, Table 1. A European-level visualization of these features is shown in Figure 3.
Resolution
@@
Missing Value
Partial features are excluded due to the high rate (above 15%) of missing values, summarized in Appendix, Table 1. For example, wind offshore from FR has a missing rate of 53.2% and is only available after 2023-08-07; solar from LV has a 75.5% missing rate and is only available after 2024-04-07; load from SK has a missing rate of 16.8% and is no longer available after 2024-07-01. For features with low missing rates (around 1%), such as load from EE and SI, we fill them using linear interpolation.
Figure 4: Structure of PriceFM. The input features $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ are passed into a fusion block to learn regional representation $\mathbf{F}_{r_{\mathrm{in}}}$. These regional representations are then stacked to form the spatial representation $\mathbf{S}$. Next, $\mathbf{S}$ is passed to the graph block to produce the representation $\mathbf{U}_{r_{\mathrm{out}}}$, which incorporates prior graph knowledge. Finally, these prior-knowledge-injected representations are fed into hierarchical quantile heads to produce multi-region, multi-timestep, and multi-quantile predictions.
Model
PriceFM, illustrated in Figure 4, aims to produce joint forecasts for multi-region, multi-timestep, and multi-quantile settings. The inputs $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{2}}$ for input region $r_{\mathrm{in}}$} are described in Section Preliminary. The inclusion of price and exogenous features introduces heterogeneity along the temporal and feature dimensions.
Backbone
We design a Fusion Block to enable effective fusion of heterogeneous multivariate input features and produce a spatial representation, and a Graph Block to incorporate prior graph knowledge to enforce spatial decay.
Fusion Block
We first align the temporal dimensions by projecting the price feature to length $L+\mathcal{T}$ via a dense layer with linear activation, followed by a transformation into a hidden space of dimension $h$ via $k$ dense layers:
@@
$\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h},$ (2)
${\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h}.$ (3)
@@
This fusion process enables feature aggregation across input regions with varying temporal and feature dimensions.
Graph Block
We first construct graph distance by performing a breadth-first search (BFS) traversal on the cross-border grid topology, which records the prior graph knowledge. For a given output region $r_{\mathrm{out}}\in\mathcal{R}$, we define the graph distance $d(r_{\mathrm{in}},r_{\mathrm{out}})$ as the minimal number of transmission hops from each input region $r_{\mathrm{in}}$ to out region $r_{\mathrm{out}}$, based on direct or indirect physical connectivity:
@@
where $r_{\mathrm{in}}\sim r_{\mathrm{out}}$ denotes that two regions are directly connected by a transmission line. For example, let $r_{\mathrm{out}}=\text{AT}$. Then $d(\text{AT},\text{AT})=0$. The region HU is directly connected to AT, thus $d(\text{HU},\text{AT})=1$. SK is not directly connected to AT but directly connected to HU, yielding $d(\text{SK},\text{AT})=2$.
Next, we introduce a graph decay mechanism to incorporate spatial prior knowledge. Intuitively, input regions that are topologically closer to the output region $r_{\mathrm{out}}$ are expected to exert a stronger influence than more distant ones. To formalize this intuition, we design a decay function that modulates the contribution of each neighboring region based on its graph distance, yielding a decay weight:
@@
where $w(d;c;D)\in[0,1]$ is the decay weight, $d$ is the graph distance, $c\in[-1,1]$ is the curvature parameter, and $D$ is the maximum reachable distance from the output region $r_{\mathrm{out}}$. As shown in Figure 4, for $c>0$ (Regime I), the decay weight drops off sharply with distance, meaning distant regions contribute very little. For $c=0$ (Regime II), the function reduces to linear decay, decreasing proportionally with graph distance. For $c<0$ (Regime III), the decay weight decreases more gradually, preserving the influence of neighbors. Using this decay function, the decay mask is constructed as:
@@
where $\mathbf{W}_{r_{\mathrm{out}}}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is the decay mask vector for $r_{\mathrm{out}}$, and $\mathbf{1}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is a vector of ones.
This operation acts as spatial regularization and eliminates the need for an exhaustive learning process to determine spatial weights, as required in methods such as attention.
Head
We design a multi-region, multi-timestep, and multi-quantile head, where the model produces joint probabilistic forecasts. To prevent quantile crossing issue [^1] Quantile crossing refers to the phenomenon where upper quantile predictions (e.g., 90%) fall below lower quantiles (e.g., 10%), violating the monotonicity of the quantile function. (Chernozhukov, Fernández-Val, and Galichon 2010)., we adopt a hierarchical quantile head (Yu et al. 2025). In detail, the median quantile ($\tau=0.5$) price trajectory, which represents the full set of timesteps $\mathcal{T}$, is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:

### v2

@@ Abstract
Electricity price forecasting in Europe presents unique challenges due to the continent’s increasingly integrated and physically interconnected power market. While recent advances in deep learning and foundation models have led to substantial improvements in general time series forecasting, most existing approaches fail to capture the complex spatial interdependencies and uncertainty inherent in electricity markets. In this paper, we address these limitations by introducing a comprehensive and up-to-date dataset across 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01. Building on this groundwork, we propose PriceFM, a spatiotemporal foundation model that integrates graph-based inductive biases to capture spatial interdependencies across interconnected electricity markets. The model is designed for multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting. Extensive experiments and ablation studies confirm the model’s effectiveness, consistently outperforming competitive baselines and highlighting the importance of spatial context in electricity markets. The project page is at: https://runyao-yu.github.io/PriceFM/.
1 Introduction
@@
The European electricity market is physically interconnected through a network of cross-border transmission lines, enabling the exchange of electricity between regions and optimizing the social welfare at the European level Lago et al. (2018). However, physical constraints, such as limited transmission capacity, can restrict electricity flow between regions and lead to zonal price differences Finck (2021), illustrated in Figure 1. These price disparities highlight the spatial nature of electricity price formation. Recent studies show that electricity price dynamics are strongly influenced by spatial interdependencies and cannot be accurately captured using region-specific models Do et al. (2024). Therefore, explicitly modeling the spatial structure of the European electricity market is essential for producing accurate price forecasts.
@@
Figure 1: Spatial distribution of electricity price and number of neighboring regions. (a) Electricity prices for 38 European regions averaged from 2022-01-01 to 2025-01-01. A significant zonal price difference is observed between north and south regions. (b) Number of neighboring regions that are directly connected to certain region via transmission lines. For example, France (FR) and Portugal (PT) are directly connected to Spain (ES), thus the number of neighboring regions for ES is 2. The mean value across all regions is 3.4.
@@
Most existing studies on electricity price forecasting do not explicitly model the spatial structure and focus on a single-region market, particularly Germany Muniain & Ziel (2020); Maciejowska et al. (2021); Kitsatoglou et al. (2024), as the German market is one of the largest markets in Europe. Other studies explore forecasting methods for markets such as Denmark, Finland, and Spain, also using region-specific models Ziel & Weron (2018); Gianfreda et al. (2020); Loizidis et al. (2024). More recent works explicitly model the spatial nature of the electricity price. For instance, a Graph Convolutional Network (GCN) is applied to capture spatial interdependencies in the Nordic markets, such as Norway, Sweden, and Finland Yang et al. (2024). Moreover, an attention-based variant is developed to predict prices in certain European markets such as Austria, Germany, and Hungary Meng et al. (2024). However, these models cover only subsets of Europe and primarily produce point forecasts, failing to capture the uncertainty inherent in electricity prices.
@@
Uncertainty modeling in electricity markets is critical, as the electricity price is strongly influenced by intermittent renewable generation and fluctuating demand Lago et al. (2021). Consequently, price forecasting should extend beyond traditional pointwise forecasting to explicitly quantify uncertainties, especially for applications involving risk-sensitive applications such as energy trading and operational planning Ziel & Steinert (2018). A comprehensive survey summarizes various probabilistic forecasting approaches based on quantile regression methods Lago et al. (2021). However, these existing methods often focus on single-region markets, thereby neglecting the rich spatial information.
@@
In recent years, foundation models for time series forecasting have achieved remarkable success across diverse domains, demonstrating strong generalization capabilities by capturing complex data patterns Zhou et al. (2022); Liu et al. (2023); Nie et al. (2023); Wu et al. (2023); Wang et al. (2024), which makes them attractive candidates for electricity price forecasting. In contrast to the conventional notion of foundation models based on pretraining, these models derive their foundation from a generic architecture applicable across domains, and are trained from scratch on the target data Liang et al. (2024). However, existing models are primarily designed for general univariate or multivariate time series tasks and are not explicitly tailored to electricity markets. Moreover, some foundation models only provide pointwise forecasts and thus fail to capture the uncertainty essential for risk-aware trading decisions. Therefore, addressing the spatial interdependencies and uncertainty in electricity price forecasting necessitates a tailored foundation model.
@@
To support the development of foundation models for electricity price forecasting, there is a pressing need for high-quality, large-scale, and up-to-date datasets that reflect the spatiotemporal complexity of integrated European markets. However, existing datasets are often fragmented in structure, cover only short time periods, are outdated, or focus on individual regions Lago et al. (2021). This lack of standardized data poses a significant barrier to training and evaluating foundation models.
In this paper, we introduce a comprehensive and up-to-date dataset and propose PriceFM, a foundation model that incorporates graph-based inductive biases to generate probabilistic forecasts. Similar to other time-series foundation models, PriceFM adopts a generic architecture that can be trained from scratch and applied across various markets. Our contributions are as follows:
@@
  • We introduce and release a comprehensive, up-to-date dataset. To the best of our knowledge, this is the largest and most diverse open dataset for European electricity markets, comprising day-ahead electricity prices, day-ahead forecasts of load, solar, and wind power generation (onshore and offshore), covering 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01.
  • We propose and release the PriceFM, a novel forecasting framework that integrates prior graph knowledge derived from the spatial topology of the European electricity market. PriceFM supports joint multi-region, multi-timestep, and multi-quantile forecasting.
  • We conduct experiments to evaluate the model’s performance against multiple baselines, and assess the impact of design choices through ablation studies, thereby providing both quantitative evidence of overall performance and insights into optimal configurations.
2 Preliminary
The forecasting target is $\mathcal{T}=24$ hourly prices for the delivery day $\mathcal{D}+1$, using data available before gate closure, typically around midday on day $\mathcal{D}$. After midday on $\mathcal{D}$, the electricity prices for $\mathcal{D}+1$ are published and known. We employ a backward-looking window of size L{L} (e.g. L=24{L}=24 corresponds to 24 hours from $\mathcal{D}$), for known electricity prices, denoted as $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$. We also include forward-looking exogenous features, such as day-ahead forecasts of load, solar, and wind (onshore and offshore) power generation for $\mathcal{D}+1$, denoted as $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, made on $\mathcal{D}$ before gate closure, as well as their historical values over L{L}. The forecasting setup and the choice of feature set are widely used in prior works Maciejowska (2020); Uniejewski & Weron (2021); Meng et al. (2024). Importantly, this work aims to utilize multi-region inputs to produce multi-region, multi-timestep, and multi-quantile forecasts. Therefore, the input and output are defined as:
@@ 2 Preliminary
where $r_{\mathrm{in}},r_{\mathrm{out}}\in\mathcal{R}=\{\text{AT},\dots,\text{SK}\}$ (region codes, detailed in Appendix, Table 1.), $\tau\in\mathcal{Q}=\{0.1,0.5,0.9\}$ (quantile levels), $f_{1}=1$, and $f_{2}$ varies by region.
Figure 2: European-level energy data in 2024, averaged across regions. (a) Electricity price. Price spikes sharply during the morning and evening peak, dip around midday, and shows higher volatility in the second half of 2024. (b) Forecasted load. Load exhibits a double‐peak each day with winter peaks substantially larger than summer. (c) Forecasted solar power generation. Solar is zero overnight, rises in a smooth bell curve to a strong midday maximum, then falls back to zero by dusk, and is much higher in summer. (d) Forecasted wind power generation (onshore and offshore). Wind lacks a daily pattern, fluctuates with high‐frequency spikes, and is much higher in winter.
3 Data
3.1 Spatiotemporal Coverage
Spatially, the dataset covers 24 European countries (38 regions). These regions reflect transmission zones rather than administrative boundaries. For example, DK is split into two regions: DK1 and DK2. Each is connected to different regions, resulting in distinct cross-border power flows. Temporally, the dataset spans from 2022-01-01 to 2025-01-01, providing wide temporal coverage.
3.2 Feature Set
The feature set includes day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore). For simplicity, we refer to these features as price, load, solar, and wind (onshore and offshore), respectively. The availability of features across regions is detailed in Appendix, Table 1. A European-level visualization of these features is shown in Figure 2.
3.3 Resolution
@@
3.4 Missing Value
Partial features are excluded due to the high rate (above 15%) of missing values, summarized in Appendix, Table 1. For example, wind offshore from FR has a missing rate of 53.2% and is only available after 2023-08-07; load from SK has a missing rate of 16.8% and is no longer available after 2024-07-01. The features with low missing rates (below 1%) are filled using linear interpolation.
Figure 3: Structure of PriceFM. The input features $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ are passed into a fusion block to learn regional representation $\mathbf{F}_{r_{\mathrm{in}}}$. These regional representations are then stacked to form the spatial representation $\mathbf{S}$. Next, $\mathbf{S}$ is passed to the graph block to produce the spatial representation $\mathbf{U}_{r_{\mathrm{out}}}$. Finally, $\mathbf{U}_{r_{\mathrm{out}}}$ is fed into hierarchical quantile heads to produce joint forecasts.
4 Model
PriceFM, illustrated in Figure 3, aims to produce multi-region, multi-timestep, and multi-quantile forecasts. The inputs $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{2}}$ are described in Section Preliminary, introducing heterogeneity along the temporal and feature dimensions.
4.1 Fusion Block
We first project the temporal dimensions by transposing and projecting $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ to length $L+\mathcal{T}$ via a dense layer with linear activation, and then transpose back and project the feature dimension of both price and exogenous features into a hidden space of dimension $h$ via $k$ dense layers:
@@ 4.1 Fusion Block
$\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h},$ (2)
${\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h}.$ (3)
@@
4.2 Graph Block
We first construct graph distance by performing a breadth-first search (BFS) traversal on the cross-border grid topology, detailed in Appendix, Table 6. For a given output region $r_{\mathrm{out}}\in\mathcal{R}$, we define the graph distance $d(r_{\mathrm{in}},r_{\mathrm{out}})$ as the minimal number of transmission hops from each input region $r_{\mathrm{in}}$ to the output region $r_{\mathrm{out}}$, based on direct or indirect physical connectivity:
@@ 4.2 Graph Block
where $r_{\mathrm{in}}\sim r_{\mathrm{out}}$ denotes that two regions are directly connected by a transmission line. For example, let $r_{\mathrm{out}}=\text{AT}$. Then $d(\text{AT},\text{AT})=0$. The region HU is directly connected to AT, thus $d(\text{HU},\text{AT})=1$. SK is indirectly connected to AT via HU, yielding $d(\text{SK},\text{AT})=2$.
Next, we introduce a graph decay mechanism to inject prior graph knowledge. Intuitively, input regions that are topologically closer to the output region $r_{\mathrm{out}}$ are expected to exert a stronger influence than more distant ones. To formalize this intuition, we design a decay function that modulates the contribution of each neighboring region based on its graph distance, yielding a decay weight:
@@
where $w(d;c;D)\in[0,1]$ is the decay weight, $d$ is the graph distance, $c\in[-1,1]$ is the curvature parameter, and $D$ is the maximum reachable distance from the output region $r_{\mathrm{out}}$. As shown in Figure 3, for $c>0$ (Regime I), the decay weight drops off sharply with distance, meaning distant regions contribute very little. For $c=0$ (Regime II), the function reduces to linear decay, decreasing proportionally with graph distance. For $c<0$ (Regime III), the decay weight decreases more gradually, preserving the influence of neighbors. We construct the decay mask as:
@@
where $\mathbf{1}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is a vector of ones.
This operation acts as spatial regularization and eliminates the need for an exhaustive learning process to determine spatial weights, as required in methods such as the attention mechanism.
4.3 Head
We design a multi-region, multi-timestep, and multi-quantile head, where the model produces joint probabilistic forecasts. To prevent quantile crossing issue [^1] Quantile crossing refers to the phenomenon where upper quantile predictions (e.g., 90%) fall below lower quantiles (e.g., 10%), violating the monotonicity of the quantile function. Chernozhukov et al. (2010)., we adopt a hierarchical quantile head Yu et al. (2025). In detail, the median quantile ($\tau=0.5$) price trajectory, which represents the full set of timesteps $\mathcal{T}$, is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer: